The total number of terms in the expansion of (x+a)47 - (x-a)47 after simplification will be?
 

Grade:12

1 Answers

Arun
5 years ago
Dear Akshay
 
Total number of terms in (x+a)^47 = 48 terms
Now
In
(x+a)^47 - (x-a)^47 = all the odd terms are cancelled out.
Hence
Total terms remaining = 48 -24 = 24
 
Regards
Arun (askIITians forum expert)